Rectangle's pereimeter is 39.6 meters, the length is 120% of the width.
What is the length and width in meters?

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

2L + 2w = 39.6
L = 1.2w
---
2L + 2w = 39.6
2(1.2w) + 2w = 39.6
2.4w + 2w = 39.6
4.4w = 39.6
w = 39.6/4.4
w = 9
---
L = 1.2w
L = 1.2*9
L = 10.8
---
answer:
w = 9 m
L = 10.8 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by stanbon(75887)

Rectangle's pereimeter is 39.6 meters, the length is 120% of the width.
What is the length and width in meters?
----------------------------
Let width be "w".
The length = "1.2w".
----
Perimeter = 2(length + width)
39.6 = 2(1.2w + w)
19.8 = 2.2w
width = 9 meters
length = 1.2*9 = 10.8 meters
=======================
Cheers,
Stan H.